Q1)

Consider two events P and Q.

a. Write the general formula used to calculate the probability that either event P occurs or Q occurs or both occur.

b. How does this formula change if:

i. Events P and Q are disjoint (i.e., mutually exclusive of each other).

ii. Events P and Q are nondisjoint events that are statistically independent of each other.

iii. Events P and Q are nondisjoint events that are statistically dependent of each other.

Q2)

Rewrite the following stochastic events in mathematical notation using a suitably defined random variable:

a. The probability that the deflection at the most critical element of a certain structure exceeds 1.80 inches is 0.85.

b. Of every 100 road construction projects, 28 experience two or more fatalities.

c. Twenty percent of the time, a certain aging engineering system component breaks down at least four times during operation every year.

d. Of all college students 75% read their campus newspaper at least once a week.

@3) and Q4) in pic The following problems are from the courses textbook: Introduction to Civil Engineering Systems, Samuel Labi, Wiley, 2014 (I
Answer #1

1(a)

The general formula of probability when either of two event occur is given as

P(either A or B occurs) = P(AUB) =P(A)+ P(B)

1(b)

(i) when two events are mutually exclusive

P(A ∪ B) = P(A) + P(B).

(ii) when both events are statically independent to each other then

P (P ∩ Q) = P (P) . P (Q)

but for two disjoint event

P (P ∩ Q) = 0

(iii) when both events dependent to each other

when occurrence of one event affect the other, then it is called dependent events

P(A and B) = P(A) · P(B|A)

where P(B|A) probability of A under condition B. ( conditional probability)